# BIS Reasoning 1.0: The First Large-Scale Japanese Benchmark for Belief-Inconsistent Syllogistic Reasoning

### v4

@@ Abstract
We present BIS Reasoning 1.0, the first large-scale Japanese dataset of syllogistic reasoning problems explicitly designed to evaluate belief-inconsistent reasoning in large language models (LLMs). Unlike prior datasets such as NeuBAROCO and JFLD, which focus on general or belief-aligned reasoning, BIS Reasoning 1.0 introduces logically valid yet belief-inconsistent syllogisms to uncover reasoning biases in LLMs trained on human-aligned corpora. We benchmark state-of-the-art models—including GPT models, Claude models, and leading Japanese LLMs—revealing significant variance in performance, with GPT-4o achieving 79.54% accuracy. Our analysis identifies critical weaknesses in current LLMs when handling logically valid but belief-conflicting inputs. These findings have important implications for deploying LLMs in high-stakes domains such as law, healthcare, and scientific literature, where truth must override intuitive belief to ensure integrity and safety.
@@
1 Introduction
@@
A major concern in current LLM research is their susceptibility to human-like cognitive biases, notably the belief bias—accepting conclusions aligned with prior beliefs regardless of logical validity. This bias poses significant risks in applications requiring impartial logic. For example, Dasgupta et al. demonstrated that LLMs frequently endorse logically invalid arguments simply because their conclusions appear believable, highlighting a critical vulnerability.
@@
To address this gap, we introduce BIS Reasoning 1.0—the first Japanese dataset designed explicitly for assessing belief-inconsistent syllogistic reasoning in LLMs. BIS (Belief-Inconsistent Syllogisms) aims to evaluate whether models can uphold logical validity when correct conclusions conflict with typical beliefs or factual knowledge. Figure 1 shows a representative example from the BIS Dataset, illustrating how logically valid conclusions can defy intuitive beliefs. Specifically, our contributions include:
@@
    Comprehensive Evaluation of Leading LLMs: We benchmark state-of-the-art models—including GPT-4o, GPT-4-turbo, Claude, and prominent Japanese LLMs—under standardized conditions, providing the first systematic comparison of their performance on belief-inconsistent reasoning in Japanese.
@@
    Detailed Analysis of Performance and Bias: Our analysis identifies significant performance gaps, highlighting that even advanced LLMs struggle disproportionately with belief-inconsistent problems. We quantify these biases, investigate specific syllogistic structures prone to errors, and examine how prompts and chain-of-thought approaches affect reasoning accuracy.
@@
2 Related Work
Evaluating the logical reasoning abilities of large language models (LLMs) has become a key research focus. Benchmarks like ReClor Yu et al. (2025) and LogiQA Liu et al. (2023) use multiple-choice logic problems derived from standardized exams to test inference beyond surface semantics. Despite advances such as chain-of-thought prompting and neuro-symbolic modeling Kojima et al. (2022), LLMs still struggle to match human performance on tasks requiring rigorous logic.
LLMs not only make logical errors but also exhibit human-like cognitive biases. One well-studied bias is belief bias—the tendency to accept conclusions that align with prior beliefs regardless of logical validity Evans et al. (1983). Studies have shown that LLMs are more accurate on belief-consistent reasoning tasks and frequently misjudge belief-inconsistent syllogisms Ando et al. (2023); Ozeki et al. (2024).
@@ 2 Related Work
In the Japanese language, recent efforts have produced datasets for logical reasoning evaluation, yet limitations remain. JFLD Morishita et al. (2024) focuses on formal deductive reasoning using artificially constructed propositions to isolate logic from world knowledge. While large in scale and diverse in structure, its use of semantically unnatural sentences prevents assessment of reasoning in realistic settings. JaNLI Yanaka and Mineshima (2021) and JAMP Sugimoto et al. (2023) explore adversarial and temporal inference respectively, but they do not test belief-inconsistent reasoning and lack syllogistic structure. NeuBAROCO Ozeki et al. (2024) closely aligns with our goals, exploring belief bias in syllogistic reasoning. However, it falls short as a comprehensive benchmark: the Japanese subset contains fewer than 800 examples for the NLI task and under 100 for the multiple-choice format, and it does not exclusively target belief-inconsistent reasoning. In contrast, BIS Reasoning 1.0 offers a focused and large-scale evaluation specifically designed to test logical robustness under belief-conflicting conditions.
@@
3 Dataset Construction
We present BIS Reasoning 1.0111https://hf.co/datasets/nguyenthanhasia/BIS_Reasoning_v1.0, a Japanese-language dataset consisting of 5,000 carefully constructed syllogistic reasoning problems. The dataset is specifically designed to test the robustness of logical inference in large language models (LLMs) under conditions of belief inconsistency, where logically valid conclusions explicitly contradict widely held commonsense beliefs. BIS Reasoning 1.0 serves as a diagnostic benchmark for probing whether LLMs can prioritize formal logic over prior-knowledge heuristics in natural language reasoning.
@@ 3 Dataset Construction
Figure 2: Combined category analysis of top raw and final semantic categories in the BIS dataset. The figure illustrates the top 5 raw categories (left), top 5 final categories (right), and the heatmap shows the relationship between these raw and final categories, highlighting category intersections and distribution patterns.
To ensure linguistic fluency and naturalness, all annotators involved in dataset construction were either native Japanese speakers or individuals with advanced Japanese proficiency. Initially, the dataset covers 46 distinct semantic categories (raw categories), ranging from concrete areas like animals, food, and weather to abstract domains such as law, logic, and emotion. These detailed raw categories were subsequently consolidated into 10 broader final categories to facilitate interpretability, ensure topic balance, and support higher-level reasoning analysis. Figure 2 summarizes the distribution and relationship between these raw and final categories, reflecting a concentration in cognitively rich areas valuable for probing belief-inconsistent reasoning in LLMs.
@@
4 Experiments
4.1 General Settings
@@
We framed the task as a binary classification using concise, instruction-based prompts in Japanese, asking models to judge logical entailment by answering “{CJK}UTF8ipxmはい” (Yes) or “{CJK}UTF8ipxmいいえ” (No). Each prompt clearly stated two premises and one conclusion, accompanied by a brief system instruction reinforcing the model’s logical reasoning role. Models were not required to provide explanations, thereby isolating pure logical inference capability from linguistic fluency or explanatory quality.
@@
Accuracy is measured as the proportion of examples for which the model outputs the correct judgment–always “{CJK}UTF8ipxmはい,” since all BIS entries are logically valid. This setup ensures that errors stem from reasoning failures, not linguistic ambiguity or semantic bias.
@@
All models in the experiments (see 4.2) were evaluated under identical zero-shot conditions, using Japanese-language prompts with consistent formatting. No fine-tuning or task-specific adaptation was applied. Evaluation was performed on the full dataset to eliminate sampling variance and enable direct comparison of out-of-the-box reasoning robustness.
4.2 Model Configuration
@@
We evaluated prominent LLMs spanning both general-purpose and Japanese-specialized categories. The general-purpose group included OpenAI’s GPT-4o and GPT-4-turbo, as well as Anthropic’s Claude-3-sonnet and Claude-3-opus. These models are designed for multilingual tasks and are optimized for general reasoning performance across domains. In contrast, the Japanese-specialized models—llm-jp-3-13b, llm-jp-3-13b-instruct3 Aizawa et al. (2024), and stockmark/stockmark-13b [^2] https://hf.co/stockmark/stockmark-13b—were trained or fine-tuned specifically on Japanese data, representing dedicated efforts to advance native Japanese LLM capabilities.
Proprietary models were accessed via official APIs, with consistent configurations including temperature set to zero to ensure deterministic outputs. Open-source models were deployed locally using standardized hardware and inference settings. All prompts were in Japanese, and all 5,000 examples in the dataset were evaluated without sampling. This setup guarantees both fairness and reproducibility across model comparisons.
4.3 Overall Model Performance
Table 1 summarizes the performance of each model on the BIS Reasoning 1.0 dataset, along with complementary results obtained on over 300 belief-inconsistent syllogistic reasoning samples from the NeuBAROCO benchmark. GPT-4o demonstrated the highest accuracy on both datasets, achieving 79.54% on BIS Reasoning 1.0 and notably excelling with 94.01% on NeuBAROCO. GPT-4-turbo and the Japanese-specialized model llm-jp-3-13b showed comparable results, each reaching around 59–60% accuracy on BIS Reasoning 1.0, but significantly lower than GPT-4o. Models specifically fine-tuned on Japanese instructions, such as llm-jp-3-13b-instruct3 and stockmark-13b, exhibited lower performance, around 40–48%, highlighting potential shortcomings in their reasoning robustness. Surprisingly, Anthropic’s Claude models underperformed drastically on the BIS Reasoning 1.0 dataset, with Claude-3-sonnet and Claude-3-opus achieving only 20.34% and 7.18%, respectively, despite relatively strong performances of 78.44% and 61.07% on NeuBAROCO.
Model BIS Reasoning 1.0 NeuBAROCO
OpenAI/GPT-4o 79.54 94.01
llm-jp/llm-jp-3-13b 59.86 67.66
OpenAI/GPT-4-turbo 59.48 67.66
llm-jp/llm-jp-3-13b-instruct3 40.90 38.32
stockmark/stockmark-13b 40.34 47.90
Claude-3-sonnet-20240229 20.34 78.44
Claude-3-opus-20240229 7.18 61.07
Table 1: Accuracy of models on the BIS Reasoning 1.0 dataset and NeuBAROCO belief-inconsistent syllogisms.
These results indicate that superior performance in belief-inconsistent reasoning is heavily influenced by a model’s underlying training approach and inherent architectural reasoning capabilities rather than mere specialization in Japanese or model size alone. Notably, GPT-4o’s ability to consistently achieve high accuracy suggests a strong robustness to belief bias, allowing it to distinguish logical validity effectively despite intuitive contradictions.
@@ 4.3 Overall Model Performance
4.4 Detailed Error Analysis for GPT-4o
To further investigate GPT-4o’s behavior on belief-inconsistent syllogisms, we conducted additional experiments on 100 cases where GPT-4o initially failed. These error samples were reassessed using several carefully designed prompts, each emphasizing different reasoning approaches, linguistic styles, or explicit instructions about the belief-inconsistent nature of the task. Figure 3 summarizes the accuracy results across these varied prompts.
Figure 3: Error Sample Accuracy by Prompt Type for GPT-4o

### v5

@@ Abstract
We present BIS Reasoning 1.0, the first large-scale Japanese dataset of syllogistic reasoning problems explicitly designed to evaluate belief-inconsistent reasoning in large language models (LLMs). Unlike prior resources such as NeuBAROCO and JFLD, which emphasize general or belief-aligned logic, BIS Reasoning 1.0 systematically introduces logically valid yet belief-inconsistent syllogisms to expose belief bias—the tendency to accept believable conclusions irrespective of validity. We benchmark a representative suite of cutting-edge models—including OpenAI GPT-5 variants, GPT-4o, Qwen, and prominent Japanese LLMs—under a uniform, zero-shot protocol. Reasoning-centric models achieve near-perfect accuracy on BIS Reasoning 1.0 (e.g., Qwen3-32B $\approx$99% and GPT-5-mini up to $\approx$99.7%), while GPT-4o attains around 80%. Earlier Japanese-specialized models underperform, often well below 60%, whereas the latest llm-jp-3.1-13b-instruct4 markedly improves to the mid-80% range. These results indicate that robustness to belief-inconsistent inputs is driven more by explicit reasoning optimization than by language specialization or scale alone. Our analysis further shows that even top-tier systems falter when logical validity conflicts with intuitive or factual beliefs, and that performance is sensitive to prompt design and inference-time reasoning effort. We discuss implications for safety-critical domains—law, healthcare, and scientific literature—where strict logical fidelity must override intuitive belief to ensure reliability.
Keywords: Belief-inconsistent reasoning, syllogistic reasoning, Japanese language models, logical benchmarking, dataset evaluation
[…]
@@
Ha-Thanh Nguyenthanks: Corresponding email: nguyenhathanh@nii.ac.jp, Hideyuki Tachibana, Chaoran Liu, Qianying Liu,
Su Myat Noe, Koichi Takeda, Sadao Kurohashi
Research and Development Center for Large Language Models, NII, Tokyo, Japan
Abstract content
1.  Introduction
@@
A major concern in current LLM research is their susceptibility to human-like cognitive biases, notably the belief bias – accepting conclusions aligned with prior beliefs regardless of logical validity. This bias poses significant risks in applications requiring impartial logic. For example, Dasgupta et al. (2022) demonstrated that LLMs frequently endorse logically invalid arguments simply because their conclusions appear believable, highlighting a critical vulnerability.
@@
To address this gap, we introduce BIS Reasoning 1.0 – the first Japanese dataset designed explicitly for assessing belief-inconsistent syllogistic reasoning in LLMs. BIS (Belief-Inconsistent Syllogisms) aims to evaluate whether models can uphold logical validity when correct conclusions conflict with typical beliefs or factual knowledge. Fig.˜1 shows a representative example from the BIS Dataset, illustrating how logically valid conclusions can defy intuitive beliefs. Specifically, our contributions include:
@@
    Comprehensive Evaluation of Leading LLMs: We benchmark state-of-the-art models – including OpenAI GPT, Claude, Qwen and prominent Japanese LLMs – under standardized conditions, providing the first systematic comparison of their performance on belief-inconsistent reasoning in Japanese.
@@
    Detailed Analysis of Performance and Bias: Our analysis identifies significant performance gaps, highlighting that even advanced LLMs struggle disproportionately with belief-inconsistent problems. We quantify these biases, investigate specific syllogistic structures prone to errors, and examine how prompts and CoT (chain-of-thought) approaches affect reasoning accuracy.
@@
2.  Related Work
Evaluating the logical reasoning abilities of large language models (LLMs) has become a key research focus. Benchmarks like ReClor Yu et al. (2025) and LogiQA Liu et al. (2023) use multiple-choice logic problems derived from standardized exams to test inference beyond surface semantics. Despite advances such as chain-of-thought prompting (Wei et al., 2022; Kojima et al., 2022) and neuro-symbolic modeling, LLMs still struggle to match human performance on tasks requiring rigorous logic.
LLMs not only make logical errors but also exhibit human-like cognitive biases. One well-studied bias is belief bias – the tendency to accept conclusions that align with prior beliefs regardless of logical validity Evans et al. (1983). Studies have shown that LLMs are more accurate on belief-consistent reasoning tasks and frequently misjudge belief-inconsistent syllogisms Ando et al. (2023); Ozeki et al. (2024).
@@ 2.  Related Work
In the Japanese language, recent efforts have produced datasets for logical reasoning evaluation, yet limitations remain. JFLD Morishita et al. (2024) focuses on formal deductive reasoning using artificially constructed propositions to isolate logic from world knowledge. While large in scale and diverse in structure, its use of semantically unnatural sentences and synthetic vocabulary prevents assessment of reasoning in realistic settings. JaNLI Yanaka and Mineshima (2021) and JAMP Sugimoto et al. (2023) explore adversarial and temporal inference respectively, but they do not test belief-inconsistent reasoning and lack syllogistic structure. NeuBAROCO Ozeki et al. (2024) closely aligns with our goals, exploring belief bias in syllogistic reasoning. However, it falls short as a comprehensive benchmark: the Japanese subset contains fewer than 800 examples for the NLI task and under 100 for the multiple-choice format, and it does not exclusively target belief-inconsistent reasoning. In contrast, BIS Reasoning 1.0 offers a focused and large-scale evaluation specifically designed to test logical robustness under belief-conflicting conditions.
@@
Figure 2: Combined category analysis of raw categories (left) and 10 final categories: animals & living things, ecosystem, human body & senses, geology, natural phenomena, models, logic & structure, arts, technology and society.
3.  Dataset Construction
We present BIS Reasoning 1.0, a Japanese-language dataset consisting of 5,000 carefully constructed syllogistic reasoning problems. The dataset is specifically designed to test the robustness of logical inference in large language models (LLMs) under conditions of belief inconsistency, where logically valid conclusions explicitly contradict widely held commonsense beliefs. BIS Reasoning 1.0 serves as a diagnostic benchmark for probing whether LLMs can prioritize formal logic over prior-knowledge heuristics in natural language reasoning.
@@ 3.  Dataset Construction
To ensure linguistic fluency and naturalness, all annotators involved in dataset construction were either native Japanese speakers or individuals with advanced Japanese proficiency. Initially, the dataset covers 46 distinct semantic categories (raw categories), ranging from concrete areas like animals, food, and weather to abstract domains such as law, logic, and emotion (left side of Fig.˜2). These detailed raw categories were subsequently consolidated into 10 broader final categories to facilitate interpretability, ensure topic balance, and support higher-level reasoning analysis (right side of Fig.˜2).
@@
4.  Experiments
Model BIS Reasoning 1.0 NeuBAROCO
GPT-5-mini (medium reasoning effort) 99.72 91.92
GPT-5-nano (medium reasoning effort) 98.84 91.32
gpt-oss-20b 98.56 89.52
GPT-4o 79.54 94.01
GPT-5-nano (minimum reasoning effort) 69.22 73.05
GPT-4-turbo 59.48 67.66
Qwen3-32B (w/o thinking) 99.58 94.01
Qwen3-32B (w/ thinking) 99.12 97.60
llm-jp-3-13b-instruct4 84.66 73.65
stockmark-13b 55.90 60.48
llm-jp-3-13b 34.78 36.52
llm-jp-3-13b-instruct3 11.06 22.46
Claude-3-sonnet-20240229 (deprecated) 20.34 78.44
Claude-3-opus-20240229 (deprecated) 07.18 61.07
Table 1: Accuracy of models on the BIS Reasoning 1.0 dataset and NeuBAROCO belief-inconsistent syllogisms. All the results are based on the ‘Basic’ prompt in Table˜2.
4.1.  General Settings
@@
We framed the task as a binary classification using concise, instruction-based prompts in Japanese, asking models to judge logical entailment by answering “Yes” or “No.” Each prompt clearly stated two premises and one conclusion, accompanied by a brief system instruction reinforcing the model’s logical reasoning role as shown in Table˜2. LLMs were not explicitly required nor prohibited to provide explanations. To derive the Yes/No labels from the LLM responses, we primarily utilized heuristic pattern matching. In cases where this approach fell short due to lengthy elucidations by the LLMs, we employed the LLM-as-a-judge strategy, using the Qwen3-32B model which has high reasoning capabilities as demonstrated in Table˜1.
@@
Accuracy is measured as the proportion of examples for which the model outputs the correct judgment – always “Yes,” since all BIS entries are logically valid. This setup ensures that errors stem from reasoning failures, not linguistic ambiguity or semantic bias.
@@
All models in the experiments (see Section˜4.1) were evaluated under identical zero-shot conditions, using Japanese-language prompts with consistent formatting. No fine-tuning or task-specific adaptation was applied. Evaluation was performed on the full dataset to eliminate sampling variance and enable direct comparison of out-of-the-box reasoning robustness.
Model Configuration
@@
We evaluated prominent LLMs spanning both general-purpose and Japanese-specialized categories. The general-purpose group included OpenAI’s GPT models (GPT-5-mini, GPT-5-nano, GPT-4o, GPT-4-turbo and gpt-oss-20b) and Alibaba’s Qwen3-32B. These models are designed for multilingual tasks and are optimized for general reasoning performance across domains. In contrast, the Japanese-specialized models – llm-jp-3-13b, llm-jp-3-13b-instruct3, llm-jp-3.1-13b-instruct4 Aizawa et al. (2024), and Stockmark’s stockmark-13b – were trained or fine-tuned specifically on Japanese data, representing dedicated efforts to advance native Japanese LLM capabilities.
All prompts were in Japanese (Basic prompt in Table˜2), and all 5,000 examples in the dataset were evaluated without sampling. This setup guarantees both fairness and reproducibility across model comparisons.
4.2.  Overall Model Performance
Table˜1 summarizes the performance of each model on the BIS Reasoning 1.0 dataset, along with complementary results obtained on over 300 belief-inconsistent syllogistic reasoning samples from the NeuBAROCO benchmark.
The strongest performances were achieved by recent reasoning-optimized models. Both GPT-5-mini and GPT-5-nano with medium (default) reasoning effort achieved near-perfect accuracy on the BIS Reasoning 1.0 dataset (99.7% and 98.8%, respectively) while maintaining above 91% accuracy on NeuBAROCO. Similarly, the open model Qwen3-32B reached above 99% accuracy with/without explicit thinking-style reasoning.
We also observed a clear trend that more recently developed models tend to exhibit stronger reasoning capabilities; for context, several deprecated early-2024 Claude models (Claude-3-sonnet-20240229 and Claude-3-opus-20240229) were included in the table to provide a broader comparison, achieving 20.3% and 7.2% accuracy on the BIS Reasoning 1.0 dataset, and 78.4% and 61.1% on NeuBAROCO, respectively.
Performance of Japanese-specific Models
The performance of Japanese-specialized models displayed substantial variation across generations. Earlier llm-jp variants and stockmark-13b lagged far behind, typically achieving between 10–60% accuracy, which underscores their limited capacity to override belief-consistent intuitions with formal logic. However, the most recent llm-jp-3.1-13b-instruct4 exhibited a marked improvement, reaching 84.66% accuracy—nearly on par with general-purpose reasoning-oriented models. This sharp rise suggests that newer iterations of Japanese LLMs are increasingly benefiting from fine-tuning strategies that explicitly emphasize reasoning alignment, rather than focusing solely on linguistic fluency or instruction adherence.
Notably, llm-jp-3-13b-instruct3 performed significantly below the 50% chance level. Such results do not merely indicate random degradation of performance, but rather a systematic tendency to favor belief-consistent conclusions even when they are logically invalid. During generation, these earlier models often became “confused,” producing extended deliberations that acknowledge inconsistencies or factual mismatches without converging on a final decision until reaching the maximum token limit. This behavior reflects a lack of internal mechanisms for reasoning prioritization and highlights the impact of instruction-tuning objectives that insufficiently penalize belief-driven heuristics.
In contrast, the emergence of llm-jp-3.1-13b-instruct4 demonstrates that reasoning-focused refinement can substantially enhance logical consistency even within Japanese-language models. Its strong performance implies that recent Japanese LLM development has begun to integrate more explicit reasoning objectives—aligning with the broader global trend toward reasoning-centric fine-tuning observed in models such as GPT-5 and Qwen3-32B. This transition underscores a maturing understanding that linguistic naturalness alone is insufficient; robust logical control mechanisms are essential for achieving true reasoning fidelity.
Reasoning Effort
GPT-4o, which previously dominated several reasoning benchmarks, now shows only moderate performance on BIS Reasoning 1.0 (79.5%), though it retains high accuracy on NeuBAROCO (94.0%). One plausible explanation is that the prompt format in our evaluation did not explicitly require models to produce elaborate reasoning chains beyond a simple “Yes” or “No” response. Since GPT-4o does not allocate an explicit reasoning effort comparable to that of newer reasoning-enabled models, its internal deliberation may have been suppressed, resulting in lower scores on BIS Reasoning 1.0. We will revisit this issue in more detail in Section˜4.3.
Regarding reasoning effort, a key insight emerges from the GPT-5-nano results obtained under different reasoning-effort settings: an API parameter that controls the model’s internal resources devoted to reasoning. When this parameter was set to medium, the model achieved 98.8% accuracy on BIS Reasoning 1.0, but the score dropped sharply to 69.2% when the effort was reduced to minimum. This finding clearly indicates that, for this dataset, substantial differences in performance arise depending on how inference-time reasoning capacity is activated and utilized.
Performance Gap Between BIS Reasoning 1.0 and NeuBAROCO
Surprisingly, Anthropic’s Claude models underperformed drastically on the BIS Reasoning 1.0 dataset, with Claude-3-sonnet and Claude-3-opus achieving only 20.34% and 7.18%, respectively, despite relatively strong performances of 78.44% and 61.07% on NeuBAROCO.
@@ 4.2.  Overall Model Performance
Figure 3: Category-wise reasoning accuracy for each LLM and final-category.
Prompt Type Prompt Description (Japanese) Prompt Description (English)
Basic 以下の三段論法を考えてください:結論は前提から論理的に導き出されますか?『はい』か『いいえ』で答えてください。[…] Consider the following syllogism: Is the conclusion logically derived from the premises? Answer "yes" or "no." […]
Focus Logic これは信念の不一致を含む可能性のある論理推論のサンプルです。前提と結論の論理的な関係に焦点を当て、結論が前提から確実に導き出されるかどうかを厳密に判断してください。回答は『はい』または『いいえ』のみでお願いします。[…] This is a sample of logical reasoning that may involve belief inconsistency. Focus strictly on the logical relationship between premises and conclusion, and determine rigorously if the conclusion follows necessarily from the premises. Answer only "yes" or "no." […]
Chain-of-Thought 以下の三段論法について、結論が前提から論理的に導き出されるかどうかを段階的に考えてください。まず、前提を分析し、次にそれらの関係性を検討し、最後に結論が論理的に妥当であるかを判断してください。思考プロセスを詳細に記述した後、最終的な回答として『はい』または『いいえ』を明確に示してください。[…] Consider the following syllogism step-by-step to determine if the conclusion logically follows from the premises. First, analyze the premises, then consider their relationship, and finally judge if the conclusion is logically valid. Provide a detailed reasoning process and clearly state your final answer as "yes" or "no." […]
Polite 以下の三段論法につきまして、結論が前提から論理的に導き出されるかどうか、ご判断いただけますでしょうか。『はい』または『いいえ』にてお答えいただけますと幸いです。[…] Could you please determine whether the conclusion logically follows from the premises in the following syllogism? I would appreciate it if you answer with either "yes" or "no." […]
Casual この三段論法、どうかな?結論、前提から合ってる?『はい』か『いいえ』で教えて。[…] What do you think about this syllogism? Does the conclusion match with the premises? Let me know with "yes" or "no." […]
Table 2: Descriptions of prompt types. “Basic” prompt was used for general evaluation, and other prompts were used for re-evaluating GPT-4o errors.
Category-wise Performance Analysis
The proposed dataset also allows for assessment by category. Fig.˜3 presents the reasoning performance across individual categories. In general, the success of belief-inconsistent reasoning does not significantly depend on the topic for most models examined. Top-tier models (Qwen3-32B, gpt-oss-20b and GPT-5 models with medium reasoning effort) simply excel at counterfactual reasoning, with no notable disparities across the various categories. In contrast, while the LLM-jp models exhibit a degree of diversity, their performance levels are generally observed to be lower.
Conversely, GPT models with lower reasoning performance (GPT-4o, GPT-4-turbo and GPT-5-nano with minimal reasoning effort) demonstrate a moderate trend in their strengths and weaknesses. These models perform the best in TEC, well in HUM, ART and NAT, yet struggle in prioritizing formal logic over common beliefs in ANI (Animals & Living Things). Although the specific reasons remain unidentified due to the proprietary nature of the GPT models, it is plausible that such trends can be attributed to the topic distribution in the training datasets utilized by the development team.
Interestingly, we can also observe moderate similarity between llm-jp-3-13B-Instruct3 and llm-jp-3.1-13B-Instruct4. Their relative strengths and weaknesses correlate in a comparable manner, which may indicate the presence of shared belief biases in their training corpora.
4.3.  Detailed Error Analysis for GPT-4o
Figure 4: Error sample accuracy by prompt type for GPT-4o.
Figure 5: Prompt type accuracy on error samples (%) – retest with English prompts.
To further investigate GPT-4o’s behavior on belief-inconsistent syllogisms, we conducted additional experiments on 100 cases where GPT-4o initially failed. These error samples were reassessed using several carefully designed prompts, each emphasizing different reasoning approaches, linguistic styles, or explicit instructions about the belief-inconsistent nature of the task. Fig.˜5 summarizes the accuracy results across these varied prompts.